OpenStudy (anonymous):

solve by elimination 2x+3y=3 and 4x+6y=6

OpenStudy (anonymous):

2x+3y=3 (1) 4x+6y=6 (2) Because (2)=2*(1) there are infinite solutions as they are the same line.

OpenStudy (anonymous):

through elimination I got 12x+18y=18 and 12x+18y=18.......they are the same so I thought this meant there is no solution, Im having a hard time figuring this out so is that right/

OpenStudy (anonymous):

If they are exactly the same, then they have infinite solutions. If only the gradient is the same then there are no solutions. (Remember must be in the form y=mx+c to check the gradient m) If they are neither of the above two, and are both straight lines, then there is one solution.